### v2

# GraphCheck: Multi-Path Fact-Checking with Entity-Relationship Graphsthanks: Code available at: https://github.com/windowh1/graphcheck

Abstract
Automated fact-checking aims to assess the truthfulness of textual claims based on relevant evidence. However, verifying complex claims that require multi-hop reasoning remains a significant challenge. We propose GraphCheck, a novel framework that transforms claims into entity-relationship graphs for structured and systematic verification. By explicitly modeling both explicit and latent entities and exploring multiple reasoning paths, GraphCheck improves verification robustness. While GraphCheck excels in complex scenarios, it may be unnecessarily elaborate for simpler claims. To address this, we introduce DP-GraphCheck, a variant that employs a lightweight strategy selector to adaptively choose between direct prompting and GraphCheck. This selective mechanism improves both accuracy and efficiency by applying the appropriate level of reasoning to each claim. Experiments on the HOVER and EX-FEVER datasets demonstrate that our approach outperforms existing methods, particularly on multi-hop claims. Moreover, the strategy selection mechanism in DP-GraphCheck generalizes well to other fact-checking pipelines, highlighting the versatility of our framework.
[…]
tcb@breakable
GraphCheck: Multi-Path Fact-Checking with Entity-Relationship Graphsthanks: Code available at: https://github.com/windowh1/graphcheck
Hyewon Jeon, Jay-Yoon Lee Seoul National University {pingpong0926,lee.jayyoon}@snu.ac.kr
1 Introduction
Automated fact-checking is a task that assesses the truthfulness of claims based on relevant evidence. With a standard pipeline that includes claim detection, evidence retrieval, and veracity assessment, automated systems enhance efficiency and accuracy in fact-checking Guo et al. (2022). However, verifying complex claims that require multi-hop reasoning remains a significant challenge. Such claims often consist of interwoven subclaims, making them difficult to verify at once. Also, relevant evidence is likely dispersed across multiple documents, complicating the retrieval process (Jiang et al., 2020; Aly et al., 2021). Another key obstacle is latent entities—references not explicitly stated in the text.
For example, in the claim “The musician, who is part of Tall Birds, is a percussionist for a band that formed in Issaquah, Washington”, the phrases “The musician” and “a band” are latent entities. While these phrases correspond to specific entities (Davey Brozowski and Modest Mouse, respectively), the claim only implies relationships to these entities without explicitly revealing them. Identifying latent entities is crucial, as they can provide pivotal information for evidence retrieval and claim verification. Also, the order of latent entity identification matters, as some entities can be inferred more easily due to stronger contextual clues, which in turn help identify others as well. Conversely, initially misidentifying a challenging entity introduces false information, hindering subsequent identification steps. Thus, latent entity identification plays an important role in fact-checking.
Recent work has examined the application of large language models (LLMs) for verifying multi-hop claims without task-specific training, relying solely on few-shot in-context prompting (Brown et al., 2020; Dmonte et al., 2024). While this setting is attractive for its scalability and generality, it introduces unique challenges. Prior approaches typically leverage LLMs to generate verification paths by identifying “check-worthy” components Guo et al. (2022), such as subclaims and question-answering steps for latent entity identification (Pan et al., 2023b, a; Wang and Shu, 2023). However, due to the inherent ambiguity of check-worthiness, verification paths often lack granularity or omit key components. Moreover, these approaches typically rely on an LLM-generated, fixed verification path that may not be optimal, ultimately limiting verification accuracy.
Figure 1: Overview of GraphCheck. Compared to prior approaches, GraphCheck offers a comprehensive yet fine-grained claim decomposition. It also systematically explores multiple paths instead of following a fixed path. The overall process consists of three steps: (1) A claim is converted into a structured entity-relation graph in which both explicit and latent entities are represented. (2) Latent entities are then identified through text infilling. Multiple infilling paths are explored via graph traversal, resulting in multiple infilled graphs. (3) For each infilled graph, every triplet is individually verified. The claim is classified as Supported if at least one path yields a graph in which all triplets are verified as Supported; otherwise, the claim is classified as Not Supported.
To address these limitations, we propose GraphCheck, in which LLMs transform text-based claims into structured entity-relationship graphs through which diverse verification paths can be generated. These graphs consist of fact triplets, each defining a relationship between entities and serving as an independently verifiable subclaim. Compared to prior approaches that extract check-worthy components without a clear structure, GraphCheck effectively performs fine-grained claim decomposition while preserving key components, enabling more comprehensive verification. Furthermore, by enabling flexible graph traversal, GraphCheck avoids reliance on a single reasoning path and allows multiple orders of latent entity identification. This flexibility increases the likelihood of capturing an optimal reasoning path, improving the robustness of verification.
While GraphCheck thoroughly verifies complex claims, it may be unnecessarily elaborate for simpler cases. Some relatively simple claims can be verified more efficiently and effectively through direct prompting (DP or Direct), where the LLM directly assesses a claim’s truthfulness based on retrieved documents. To leverage the complementary strengths of Direct and GraphCheck, we introduce DP-GraphCheck, which employs a lightweight strategy selector to adaptively choose between the two methods. Notably, this strategy selector is modular and can be easily integrated into diverse fact-checking systems beyond GraphCheck.
Experimental results on multi-hop fact-checking datasets (Jiang et al., 2020; Aly et al., 2021) demonstrate that both GraphCheck and DP-GraphCheck outperform existing methods on complex multi-hop verification tasks, highlighting the effectiveness of structuring claims as entity-relationship graphs for comprehensive and systematic fact-checking. Furthermore, we find that the strategy selector improves performance not only in GraphCheck but also when applied to other baseline fact-checking systems, underscoring its broad applicability.
2 Related Work
LLM-based Fact-Checking
Traditional fact-checking has relied on rule-based systems or supervised models trained on annotated datasets. With the rise of large language models (LLMs), the field has shifted towards more flexible, context-aware reasoning (Guo et al., 2022; Dmonte et al., 2024). LLM-based methods utilize retrieval-augmented generation (RAG) (Lewis et al., 2021; Gao et al., 2024) and in-context learning Brown et al. (2020) for veracity prediction.
For multi-hop fact-checking, where reasoning across multiple evidence pieces is required, LLM-based strategies include iterative RAG Shao et al. (2023) and claim decomposition (Chen and Shu, 2023; Zhang and Gao, 2023; Pan et al., 2023b). GraphCheck introduces a novel approach, leveraging a graph structure for claim decomposition to enable comprehensive multi-hop verification.
Latent Entity Identification
Identifying latent entities is key to improving evidence retrieval and verification accuracy. While Prior work has used question-answering (Pan et al., 2023b, a; Wang and Shu, 2023) or RAG Shao et al. (2023) to tackle this, GraphCheck constructs an entity-relationship graph where latent entities are explicitly represented as nodes and identified through text infilling Zhu et al. (2019).
Existing methods also rely on model-driven ordering for latent entity identification.  Shao et al. (2023) and  Pan et al. (2023a) adopt an iterative approach where prior outputs guide subsequent steps, forming a linear sequence.  Pan et al. (2023b) generates an entire reasoning sequence in a single LLM API call and aggregates results from multiple paths generated by additional calls. In contrast, GraphCheck systematically explores multiple identification paths within a graph generated by a single API call, ensuring efficient and robust verification.
Graph-based Fact-Checking
Graph-based models have been utilized in fact-checking to capture relationships between entities in claims and evidence.  Zhong et al. (2020) employed semantic role labeling (SRL) to construct graphs from evidence and applied graph convolutional networks (GCNs) and graph attention networks (GATs) for logical inference.  Yuan and Vlachos (2024) extracted triplets using an OpenIE model and verified them through a natural language inference (NLI) model. Unlike prior approaches, GraphCheck constructs entity-relationship graphs that explicitly represent latent entities as placeholder nodes, leveraging the in-context learning capabilities of LLMs.
3 Methodology
3.1 GraphCheck: Graph Construction
GraphCheck first transforms a claim into an entity-relationship graph. Latent entities within the claim are detected and represented as placeholders, forming distinct nodes. To construct the graph, we leverage an LLM guided by instructions that include predefined rules and few-shot examples. The key instructions are as follows:
  • Decompose the claim into fact triplets, each serving as a basic unit of information.
  • Format a triplet as ‘subject [SEP] relation [SEP] object’, where the subject and object are noun phrases, and the relation is verb phrase, forming a complete sentence.
  • Detect and represent latent entities using placeholders, e.g., (ENT1), (ENT2).
As shown in the part of Figure 1 that explains the “(1) Graph Construction” step of GraphCheck, the generated graph consists of two sections:
  • # Latent Entities: Triplets that link latent entities to their implicit references in the claim.
  • # Triples: Triplets that capture relationships between entities.
While triplets in both section serve as subclaims requiring verification, the # Latent Entities section is set apart to ensure that contextual meaning is preserved when placeholders are introduced. For instance, in the example shown in Figure 1, replacing “the musician” with a placeholder (ENT1) could result in a loss of information. To prevent this, placeholders are explicitly mapped to their corresponding references in the seperated section.
3.2 GraphCheck: Latent Entity Infilling
Latent Entity Identification
Once the entity-relationship graph is constructed, latent entities are identified sequentially through text infilling. The process begins by retrieving top-$k$ documents relevant to the target latent entity. To ensure sufficient context, a retrieval query is generated by concatenating all triplets that include the target entity and exclude any other unidentified latent entities. The retrieved documents, along with the same set of triplets, are then used as input to the text infilling model. Within this input, the target latent entity is represented by a special token. A more detailed description of the infilling process is provided in Appendix B.2.
The part of Figure 1 that explains the (2) Latent Entity Infilling step shows an example of the infilling process. In path 1, (ENT1) is identified first as the target entity. Then, (ENT2) becomes the next target entity, at which point (ENT1) has already been identified as “Randall Nieman”, providing additional contextual information.
Multiple Path Exploration
For claims containing multiple latent entities, various identification paths are explored. In the example shown in Figure 1, two possible paths—(ENT1$\rightarrow$(ENT2), (ENT2$\rightarrow$(ENT1)—are considered. To manage computational complexity, up to $\bar{P}$ paths are randomly sampled if the total number of possible sequences exceeds this limit.
Exploring various identification paths is crucial, as some paths may be more effective than others. Certain latent entities can be identified more easily due to stronger contextual clues, and identifying them first can provide additional context for subsequent identifications. For example, in Figure 1, identifying (ENT2) is easier than identifying (ENT1) because “Issaquah, Washington” provides a more salient retrieval cue than “Tall Birds”, which may introduce ambiguity between a rock band and actual birds. As a result, when (ENT2) is identified first in Path 2, both entities are correctly infilled, whereas Path 1 fails to infill (ENT1) correctly.
However, automatically finding the optimal identification path is a challenging task. To address this, our method systematically explores multiple paths rather than relying on model-driven path planning, increasing the likelihood of finding the most effective path.
3.3 GraphCheck: Fact Triplet Verification
Triplet Verification
After infilling, each triplet in the graph is independently verified. For each infilled triplet $t^{\prime}$, the top-$k$ documents are retrieved from the corpus using $t^{\prime}$ as the retrieval query. The verifier assesses the veracity of $t^{\prime}$ using two types of evidence inputs: (i) the concatenation of the top-$k$ documents, and (ii) each of the top-$k$ documents individually, resulting in a total of $k+1$ inputs. If any of these inputs yields a Supported judgment, $t^{\prime}$ is classified as Supported; otherwise, it is classified as Not Supported.
Path Verification
A latent entity identification path produces a fully infilled graph consisting of triplets $\{t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}\}$. The path is classified as Supported if all triplets are classified as Supported; otherwise, it is classified as Not Supported.
Claim Verification
Since multiple identification paths can be explored, GraphCheck performs triplet-level verification for each path independently. A claim is ultimately classified as Supported if at least one path is classified as Supported; otherwise, it is classified as Not Supported.
This classification approach accounts for potential errors in latent entity infilling. A claim classified as Not Supported in a single identification path does not necessarily indicate that the claim itself is false; rather, it may result from improperly identified latent entities. Exploring multiple identification paths increases the likelihood of correct verification, as a claim is successfully verified if at least one path accurately identifies latent entities, thereby supporting robust fact-checking.
3.4 DP-GraphCheck
Algorithm 1 DP-GraphCheck
Input: Claim $C$
Modules: Strategy Selector $\mathcal{S}$, Verifier $\mathcal{V}$, Retriever $\mathcal{R}$
Determine strategy $\sigma\leftarrow\mathcal{S}(C)$
if $\sigma=\textsc{Direct}$ then
     Retrieve top-$k$ documents $\{d_{1},\dots,d_{k}\}\leftarrow\mathcal{R}(C)$
     Concatenate docs: $d_{\text{concat}}\leftarrow d_{1}\oplus\cdots\oplus d_{k}$
     Use verifier $\mathcal{V}(C,d_{\text{concat}})$ to classify $C$
     return $\mathcal{V}(C,d_{\text{concat}})$
end if
# GraphCheck
Construct graph $\mathcal{G}_{C}=\{t_{1},t_{2},\dots,t_{n}\}$
for each path $\pi_{p}\in\{\pi_{1},\dots,\pi_{\min(P,\bar{P})}\}$ do
     Infill graph $\mathcal{G}_{C,\pi_{p}}^{\prime}=\{t_{1}^{\prime},t_{2}^{\prime},\dots,t_{n}^{\prime}\}$
     $path\_supported\leftarrow$ True
     for each infilled triplet $t_{i}^{\prime}\in\mathcal{G}_{C,\pi_{p}}^{\prime}$ do
         Retrieve top-$k$ documents $\{d_{1},\dots,d_{k}\}\leftarrow\mathcal{R}(t_{i}^{\prime})$
         Concatenate docs: $d_{\text{concat}}\leftarrow d_{1}\oplus\cdots\oplus d_{k}$
         $triplet\_supported\leftarrow$ False
         for each $d_{j}\in\{d_{\text{concat}},d_{1},\dots,d_{k}\}$ do
              if $\mathcal{V}(t_{i}^{\prime},d_{j})=\textsc{Supported}$ then
                  $triplet\_supported\leftarrow$ True
                  break
              end if
         end for
         if not $triplet\_supported$ then
              $path\_supported\leftarrow$ False
              break
         end if
     end for
     if $path\_supported$ then
         return Supported
     end if
end for
return Not Supported
GraphCheck rigorously evaluates the supportedness of a claim; a claim is classified as Supported only if the infilling results of all triplets align with the retrieved evidence. However, for relatively simple claims that do not require decomposition or latent entity identification, this approach can be unnecessarily strict and may misclassify correct claims as Not Supported.
To address this limitation, we introduce DP-GraphCheck, which improves both efficiency and accuracy. Given the claim and its top-$k$ retrieved documents (using the original claim itself as a query), a lightweight strategy selector determines whether the retrieved evidence is sufficient for assessing its veracity. If deemed sufficient, the claim is considered simple and verified via Direct. Otherwise, the claim undergoes the full GraphCheck pipeline, including claim decomposition and latent entity infilling.
This framework efficiently filters out simple claims, while maintaining thorough verification for more complex cases. The complete verification process of DP-GraphCheck is summarized in Algorithm 1.
4 Experiments
4.1 Experimental Setup
Datasets
We utilize two datasets:
  • HOVER Jiang et al. (2020) is a dataset for multi-hop fact-checking, verifying whether a claim is supported or not based on evidence dispersed across multiple Wikipedia articles (2-hop, 3-hop, 4-hop). Since the test set labels are not publicly available, we use the development set as our test set. We use the preprocessed October 2017 Wikipedia dump Yang et al. (2018) as the retrieval corpus.
  • EX-FEVER Aly et al. (2021) is another multi-hop fact-checking dataset, where evidence is scattered across multiple Wikipedia articles (2-hop, 3-hop). Unlike HOVER, which has only two labels, EX-FEVER introduces an additional ‘Not Enough Information (NEI)’ label. To align with the HOVER dataset, we exclude samples labeled as NEI. Following the dataset authors’ guidance, we use the same October 2017 Wikipedia dump as HOVER for the retrieval corpus.
Implementation Details
We employ FLAN-T5-XL Chung et al. (2022) for text infilling, fact triplet verification, strategy selection, and Direct, applying task-specific prompts (Appendix B). This setup aligns with our baseline, ProgramFC Pan et al. (2023b), ensuring a fair comparison.
For retrieval, we employ BM25 Robertson and Zaragoza (2009), maintaining consistency with ProgramFC as well. Our primary focus is the open-book setting, where the verification is conducted based on the top-$k$ retrieved documents ($k{=}10$). We also evaluate performance in the open-book + gold setting, where the claim’s gold document set is merged with the top retrieved documents to form a set of $k$ documents.
For entity-relationship graph construction, we employ gpt-4o (gpt-4o-2024-08-06). The few-shot examples used in the graph construction prompt are manually annotated using 10 instances randomly sampled from the HOVER training set (Appendix B.1).
The path limit $\bar{P}$, which defines the maximum number of exploration paths, is set to 5. Additionally, $\bar{P}{=}1$ is also tested to assess the impact of multiple path exploration.
Baselines
We compare our approach against several fact-checking frameworks that rely on in-context learning:
  • ProgramFC Pan et al. (2023b) is a fact-checking framework that decomposes complex claims into Python-like reasoning programs, outlining step-by-step actions such as question answering and subclaim verification. We generally follow the original experimental setup; however, since ProgramFC originally used Codex Chen et al. (2021), we re-run the experiments using gpt-4o to ensure comparability. We evaluate two cases, $N{=}1$ and $N{=}5$, where $N$ represents the number of OpenAI API calls, with each call expected to generate a distinct reasoning program.
  • FOLK Wang and Shu (2023) decomposes claims into First-Order Logic (FOL) clauses and question-answering sets required for fact-checking. Originally, FOLK used gpt-3.5 (text-davinci-003) for decomposing and SerpAPI for evidence retrieval and question answering. To align with our setup, we adapt FOLK to use gpt-4o and replace its question-answering module with FLAN-T5-XL, which generates answers based on Wikipedia articles retrieved via BM25.
  • FLAN-T5-XL Chung et al. (2022), a model with 3 billion parameters, has shown strong generalization and reasoning capabilities across various tasks. As it serves as the Direct component in DP-GraphCheck, we also report its standalone performance.
Additionally, we apply the strategy selector to ProgramFC and FOLK, to evaluate its generalizability.
Methods Open-book Open-book + Gold
HOVER EX-FEVER HOVER EX-FEVER
2-hop 3-hop 4-hop 2-hop 3-hop 2-hop 3-hop 4-hop 2-hop 3-hop
Direct (DP) 72.56 61.70 59.57 81.03 73.02 76.03 67.18 61.26 87.77 81.82
ProgramFC ($N{=}1$) 70.04 61.33 59.00 77.55 71.50 71.52 64.74 63.99 83.82 78.52
DP-ProgramFC ($N{=}1$) 70.79 62.75 60.61 79.46 74.62 71.19 66.04 65.31 83.42 78.93
ProgramFC ($N{=}5$) 70.29 61.82 60.19 78.31 72.16 70.73 65.50 63.79 84.09 79.69
DP-ProgramFC ($N{=}5$) 70.71 63.39 61.56 79.78 75.30 70.52 66.23 64.86 83.69 79.80
FOLK 65.13 59.63 56.10 72.32 63.85 69.96 66.23 65.65 80.13 75.98
DP-FOLK 70.96 63.53 58.36 80.64 73.68 72.48 69.39 68.59 84.99 80.66
GraphCheck ($\bar{P}{=}1$) 73.05 64.87 59.19 75.71 65.02 78.18 70.68 67.70 83.06 75.86
DP-GraphCheck ($\bar{P}{=}1$) 76.29 67.36 62.35 81.12 74.56 77.25 73.00 71.95 85.78 82.87
GraphCheck ($\bar{P}{=}5$) 74.12 67.71 64.79 76.56 69.94 78.59 73.78 72.55 83.64 80.16
DP-GraphCheck ($\bar{P}{=}5$) 76.29 68.70 66.64 81.12 76.02 76.96 73.34 73.63 85.69 82.73
Table 1: Macro-F1 scores under open-book and open-book + gold settings. The best result in each column is highlighted in bold, and the second-best result is underlined. The ‘DP-’ notation indicates that the strategy selector is used to combine Direct with the corresponding method.
4.2 Main Result
Table 1 summarizes the Macro-F1 scores for both the open-book and open-book + gold settings. The ‘DP-’ notation indicates that the strategy selector is used to combine Direct with the corresponding method.
DP-GraphCheck
In nearly all configurations—including the HOVER and EX-FEVER dataset, under both open-book and open-book + gold settings—DP-GraphCheck consistently achieves either the best or second-best Macro-F1 score. This demonstrates the strength of our framework, which combines the best of both worlds: it maintains the efficiency and precision of direct prompting for simple claims, while leveraging the structured reasoning of GraphCheck for more complex ones. As a result, DP-GraphCheck achieves the strongest overall performance.
An exception occurs in the EX-FEVER open-book + gold 2-hop setting, where the Direct method achieves the highest score. This can be attributed to EX-FEVER’s extractive nature, where claims closely match phrases in the gold documents with minimal rephrasing or abstraction (see Appendix A for a specific example). In such cases, Direct, which preserves the claim’s original form, may yield strong results. However, this advantage may not generalize to real-world fact-checking scenarios, where claims often differ in wording from the supporting evidence. In these realistic settings, more structured reasoning becomes essential.
GraphCheck
In the HOVER dataset, under both the open-book and open-book + gold settings, GraphCheck with $\bar{P}{=}5$ outperforms all baselines, even without the strategy selector. This highlights the effectiveness of graph-based structured verification in settings that require deeper reasoning and decomposition.
In the EX-FEVER dataset, GraphCheck occasionally underperforms compared to baseline methods. Considering the extractive nature of EX-FEVER, these results may be attributed to GraphCheck’s enforced fine-grained decomposition, in which the claim is broken down into a set of entity-relation triplets, each corresponding to a subclaim. While such structured decomposition is beneficial for complex reasoning, it may be unnecessary for extractive-style claims that can be verified holistically. However, when combined with Direct (i.e., DP-GraphCheck), GraphCheck consistently recovers strong performance.
Multi-path Verification
We also observe that GraphCheck with $\bar{P}{=}5$ generally outperforms its $\bar{P}{=}1$ counterpart, and this improvement frequently carries over to the DP variants. The performance gap becomes more pronounced as the hop count increases. For example, in the HOVER open-book setting, GraphCheck ($\bar{P}{=}5$) surpasses GraphCheck ($\bar{P}{=}1$) by 1.07, 2.84, and 5.60 points in 2-hop, 3-hop, and 4-hop claims, respectively. These results indicate that multi-path exploration becomes increasingly beneficial for more complex claims.
While ProgramFC also attempts multi-path verification by generating $N$ distinct verification programs via $N$ independent LLM API calls, the performance gap between ProgramFC ($N{=}1$) and ProgramFC ($N{=}5$) remains relatively small. This suggests that the reasoning paths generated by each call may lack sufficient diversity. Moreover, this method incurs high computational cost due to repeated LLM usage, even for claims that may not require extensive reasoning.
In contrast, GraphCheck constructs a single entity-relation graph through one LLM call, from which diverse verification paths can be explored. This design naturally promotes reasoning diversity without additional LLM overhead.
Effectiveness of Strategy Selector
The strategy selector is effective across all baselines: both DP-ProgramFC and DP-FOLK consistently outperform their original versions. These results show that the strategy selector not only strengthens our own framework, but also serves as a general enhancement applicable to existing fact-checking methods.
4.3 Ablation Study
Breakdown of DP-GraphCheck
Table 2 presents a detailed breakdown of DP-GraphCheck’s performance on the HOVER dataset under the open-book setting. It reports results for the entire dataset as well as for two subsets of claims, each assigned to either Direct or GraphCheck by the strategy selector. For each group, the table shows the proportion of claims, the recall of retrieval results using the original claim as the query (i.e., the evidence provided to both the strategy selector and Direct), and the verification accuracy of both Direct and GraphCheck.
We observe that the strategy selector assigns an increasing proportion of claims to GraphCheck as hop count increases: 59.68% of 2-hop claims are handled by GraphCheck, compared to 88.35% for 4-hop claims. This trend indicates that the strategy selector is capable of discerning claim difficulty and assigning complex cases to a more systematic verification method.
Moreover, claims verified via Direct exhibit consistently higher retrieval recall than those verified via GraphCheck. This suggests that the strategy selector effectively assesses whether retrieved evidence is sufficient to support Direct verification, thereby avoiding unnecessary decomposition.
In terms of accuracy, within the group of claims verified via Direct, the accuracy of the Direct consistently exceeds that of GraphCheck across all hop levels, demonstrating the effectiveness of direct prompting for simpler claims. Conversely, for claims verified via GraphCheck, the GraphCheck method outperforms Direct, particularly on 3-hop and 4-hop claims, highlighting the importance of structured reasoning in complex scenarios.
Overall, These results indicate that DP-GraphCheck successfully combines the strengths of Direct and GraphCheck through the strategy selector, achieving not only greater efficiency but also improved overall performance.
Group Metric 2-hop 3-hop 4-hop
Total Recall@10 73.18 51.34 36.43
Accuracy (Direct) 72.56 62.02 59.58
Accuracy (GraphCheck) 74.60 68.12 66.79
Verified via Direct % of Claims 40.32% 23.22% 11.65%
Recall@10 84.47 64.32 55.79
Accuracy (Direct) 74.01 71.13 72.73
Accuracy (GraphCheck) 69.82 68.54 64.46
Verified via GraphCheck % of Claims 59.68% 76.78% 88.35%
Recall@10 37.71 29.76 21.66
Accuracy (Direct) 71.58 59.26 57.84
Accuracy (GraphCheck) 71.48 67.99 67.10
Table 2: Breakdown of DP-GraphCheck performance on the HOVER dataset under the open-book setting. Results are grouped by the strategy selector’s assignment: total samples, samples verified via Direct, and those verified via GraphCheck. Bold values indicate the accuracy of the verification method actually used for each group.
Impact of Graph Construction Model
A core component of GraphCheck is its graph construction step, which requires decomposing a claim into fact triplets and detecting latent entities simultaneously. To assess the generalizability of GraphCheck across different language models beyond our default choice (gpt-4o), we experiment with several LLMs using identical prompts.
As shown in Table 3, smaller models such as gpt-3.5-turbo and Qwen2.5-72B-Instruct yield relatively lower performance in GraphCheck compared to gpt-4o, indicating that model capacity can influence the quality of graph construction. However, Qwen2.5-72B-Instruct, despite its relatively modest size, demonstrates surprisingly competitive results. Moreover, the performance gap between Qwen2.5-72B-Instruct and gpt-4o narrows considerably in DP-GraphCheck. For 2-hop claims, Qwen2.5-72B-Instruct achieves an even higher score (77.00 vs. 76.29), suggesting that the strategy selector effectively compensates for imperfect graph construction by leveraging direct prompting when appropriate.
In the case of claude-3-7-sonnet, the model demonstrates strong performance in both GraphCheck and DP-GraphCheck method, outperforming gpt-4o on 2-hop and 3-hop claims.
These findings suggest that the strong performance of GraphCheck and DP-GraphCheck with gpt-4o can be reproduced with other language models as well. While graph construction quality may vary depending on the underlying model—particularly for smaller LLMs—the DP mechanism helps compensate for suboptimal graphs by leveraging Direct when it is more appropriate. This demonstrates that GraphCheck and DP-GraphCheck are model-agnostic frameworks that generalize well across a range of LLM backbones.
Graph Construction Model Method 2-hop 3-hop 4-hop
gpt-4o-2024-08-06 (Default) GraphCheck 74.12 67.71 64.79
DP-GraphCheck 76.29 68.70 66.64
gpt-3.5-turbo GraphCheck 70.06 59.93 59.35
DP-GraphCheck 74.42 63.68 62.00
Qwen2.5-72B-Instruct GraphCheck 73.04 62.33 62.74
DP-GraphCheck 77.00 66.49 65.55
claude-3-7-sonnet-20250219 GraphCheck 76.39 68.53 63.65
DP-GraphCheck 76.45 69.74 65.50
Table 3: Macro-F1 scores of GraphCheck and DP-GraphCheck on the HOVER dataset (open-book setting) using different graph construction models.
Various Document-level Settings
We compare three document-level settings for verifying claims based on the top-$k$ retrieved documents: (1) concat, where the claim is verified once over the concatenation of all documents; (2) each, where the claim is evaluated against each document individually and considered Supported if any yields a Supported judgment; and (3) concat+each, which combines both approaches and returns Supported if either the concatenated context or any individual document supports the claim.
In DP-GraphCheck, Direct adopts the concat setting, while GraphCheck employs the concat+each setting. As shown in Table 4, these choices align well with the nature of input claim each method receives.
For Direct, the input is the original multi-hop claim, which typically integrates information distributed across multiple documents. Verifying the claim on each document in isolation may fail to capture this broader context, likely degrading performance. As a result, evaluating claims over the concatenated evidence (concat) shows the strongest performance.
In contrast, GraphCheck verifies fine-grained subclaims, each represented as an entity-relation triplet. These subclaims often correspond to atomic facts localized within individual documents, making the each setting more appropriate. However, some triplets still require cross-document context, which makes concat+each the most effective setting for GraphCheck.
These results show the importance of aligning the document-level setting with the nature of the input claim—whether it is a original multi-hop claim or a decomposed triplet—to achieve optimal verification performance.
Method Document-level Setting 2-hop 3-hop 4-hop
Direct concat 72.56 61.70 59.57
each 65.55 59.17 57.70
concat+each 67.90 59.92 58.40
GraphCheck concat 69.39 62.64 60.07
each 74.79 66.19 62.36
concat+each 74.12 67.71 64.79
Table 4: Macro-F1 scores of Direct and GraphCheck under different document-level settings on the HOVER dataset (open-book setting). Bold values indicate the actual document-level setting used by each method.
5 Conclusion
We introduce GraphCheck, a novel framework for automated fact-checking based on entity-relationship graphs, specifically designed to handle multi-hop reasoning. By converting claims into structured entity-relationship graphs and exploring multiple identification paths, our approach enables comprehensive and systematic verification. Additionally, we proposed DP-GraphCheck, which leverages a lightweight strategy selector to adaptively choose between simple direct prompting and more systematic graph-based verification, thereby enhancing both efficiency and robustness.
Our experiments on the HOVER and EX-FEVER datasets demonstrate that DP-GraphCheck consistently achieves superior performance across multiple settings. Furthermore, the strategy selector generalizes well to other fact-checking methods, making it a practical and adaptable component for broader use. Ablation studies further validate the effectiveness of each component, including multi-path traversal and document-level verification strategies. Overall, our findings highlight GraphCheck and DP-GraphCheck as strong and extensible frameworks for multi-hop claim verification.
Limitations
Despite its advantages, our framework has some limitations. First, the construction of entity-relationship graphs can be error-prone, which may propagate to the verification stage. Common issues include misclassifying non-latent entities as latent and failing to detect actual latent entities. Although this work focuses on a few-shot prompting setting, we believe that incorporating training could improve the quality of graph construction.
Second, our framework currently operates solely over a textual knowledge base. Given that the verification process is grounded in a structured graph representation, future extensions could explore applying GraphCheck in settings where the evidence is sourced from structured knowledge bases such as knowledge graphs.
Appendix A An Example from the EX-FEVER Dataset
EX-FEVER is highly extractive compared to the HOVER dataset. An example from EX-FEVER is shown in Figure 2. As illustrated in the example, claims in EX-FEVER closely match phrases in the gold documents with minimal rephrasing or abstraction.
Figure 2: An example from the EX-FEVER dataset.
Appendix B Prompts
B.1 Graph Construction
The prompt in Table LABEL:tab:prompt_graph_construction is used in GraphCheck to convert a textual claim into a structured entity-relationship graph. It consists of an instruction segment and 10 illustrative examples manually created using claims from the HOVER training set.
Table 5: Prompt used for graph construction.
We are conducting fact-checking on multi-hop claims. To facilitate this process, we need to decompose each claim into triples for more granular and accurate fact-checking. Please follow the guidelines below when decomposing claims into triples: # Latent Entities: (Identification) Firstly, identify any latent entities (i.e., implicit references not directly mentioned in the claim) that need to be clarified for accurate fact-checking. (Definition) Define these identified latent entities in triple format, using placeholders like (ENT1), (ENT2), etc. # Triples: (Basic Information Unit) Decompose the claim into triples, ensuring you reach the most fundamental verifiable information while preserving the original meaning. Be careful not to lose important information during decomposition. (Triple Structure) Each triple should follow this format: ‘subject [SEP] relation [SEP] object’. Both the subject and object should be noun phrases, while the relation should be a verb or verb phrase, forming a complete sentence. (Prepositional Phrases) In exceptional cases where a prepositional phrase modifies the entire triple (rather than just the subject or object) and splitting it into another triple would alter the meaning of the claim, do not divide it. Instead, append it to the end of the triple: ‘subject [SEP] relation [SEP] object [PREP] preposition phrase’. (Pronoun Resolution) Replace any pronouns with the corresponding entities to ensure that each triple is self-contained and independent of external context. (Entity Consistency) Use the exact same string to represent entities (i.e., the ‘subject’ or ‘object’) whenever they refer to the same entity across different triples. # Claim:
The fairy Queen Mab orginated with William Shakespeare.
# Latent Entities:
# Triples:
The fairy Queen Mab [SEP] originated with [SEP] William Shakespeare
# Claim:
Giacomo Benvenuti and Claudio Monteverdi share the profession of Italian composer.
# Latent Entities:
# Triples:
Giacomo Benvenuti [SEP] is [SEP] Italian composer
Claudio Monteverdi [SEP] is [SEP] Italian composer
# Claim:
Ross Pople worked with the English composer Michael Tippett, who is known for his opera “The Midsummer Marriage”.
# Latent Entities:
# Triples:
Ross Pople [SEP] worked with [SEP] the English composer Michael Tippett
The English composer Michael Tippett [SEP] is known for [SEP] the opera “The Midsummer Marriage”
# Claim:
Mark Geragos was involved in the scandal that took place in the 1990s.
# Latent Entities:
(ENT1) [SEP] is [SEP] a scandal
# Triples:
Mark Geragos [SEP] was involved in [SEP] (ENT1)
(ENT1) [SEP] took place in [SEP] the 1990s
# Claim:
Where is the airline company that operated United Express Flight 3411 on April 9, 2017 on behalf of United Express is headquartered in Indianapolis, Indiana.
# Latent Entities:
(ENT1) [SEP] is [SEP] an airline company
# Triples:
(ENT1) [SEP] operated [SEP] United Express Flight 3411 [PREP] on April 9, 2017 on behalf of United Express
(ENT1) [SEP] is headquartered in [SEP] Indianapolis, Indiana
# Claim:
The Skatoony has reruns on Teletoon in Canada and was shown between midnight and 6:00 on the network that launched 24 April 2006, the same day as rival Nick Jr. Too.
# Latent Entities:
(ENT1) [SEP] is [SEP] a network
# Triples:
Skatoony [SEP] has reruns on [SEP] Teletoon
Teletoon [SEP] is located in [SEP] Canada
Skatoony [SEP] was shown on [SEP] (ENT1) [PREP] between midnight and 6:00
(ENT1) [SEP] launched on [SEP] 24 April 2006
Nick Jr. Too [SEP] launched on [SEP] 24 April 2006
# Claim:
Danny Shirley is older than Kevin Parker.
# Latent Entities:
(ENT1) [SEP] is [SEP] a date
(ENT2) [SEP] is [SEP] a date
# Triples:
Danny Shirley [SEP] was born on [SEP] (ENT1)
Kevin Parker [SEP] was born on [SEP] (ENT2)
(ENT1) [SEP] is before [SEP] (ENT2)
# Claim:
The founder of this Canadian owned, American manufacturer of business jets for civilian and military did not develop the 8-track portable tape system.
# Latent Entities:
(ENT1) [SEP] is [SEP] an individual
(ENT2) [SEP] is [SEP] an American manufacturer
# Triples:
(ENT1) [SEP] founded [SEP] (ENT2)
(ENT2) [SEP] is owned by [SEP] Canadian
(ENT2) [SEP] made [SEP] business jets for civilian and military
(ENT1) [SEP] did not develop [SEP] 8-track portable tape system
# Claim:
The Dutch man who along with Dennis Bergkamp was acquired in the 1993-94 Inter Milan season, manages Cruyff Football together with the footballer who is also currently manager of Tel Aviv team.
# Latent Entities:
(ENT1) [SEP] is [SEP] a Dutch man
(ENT2) [SEP] is [SEP] a footballer
# Triples:
(ENT1) [SEP] was acquired in [SEP] the 1993-94 Inter Milan season [PREP] along with Dennis Bergkamp
(ENT1) [SEP] manages [SEP] Cruyff Football [PREP] together with (ENT2)
(ENT2) [SEP] currently manages [SEP] Tel Aviv team
# Claim:
An actor starred in the 2007 film based on a former FBI agent. That agent was Robert Philip Hanssen. The actor starred in the 2005 Capitol film Chaos.
# Latent Entities:
(ENT1) [SEP] is [SEP] an actor
(ENT2) [SEP] is [SEP] a 2007 film
# Triples:
(ENT1) [SEP] starred in [SEP] (ENT2)
(ENT2) [SEP] is based on [SEP] Robert Philip Hanssen
Robert Philip Hanssen [SEP] is [SEP] a former FBI agent
(ENT1) [SEP] starred in [SEP] the 2005 Capitol film Chaos
# Claim:
{claim}
B.2 Latent Entity Infilling
The prompt in Table 6 is used in the latent entity infilling step of GraphCheck to identify the target latent entity.
{context}
Based on the above information, fill in the blank with the correct entity: {infilling_query}
Answer:
Table 6: Prompt used for latent entity infilling.
The {context} refers to the concatenation of top-$k$ documents retrieved using a custom retrieval query. This retrieval query is constructed by concatenating all triplets in the # Triples section that include the target latent entity to be infilled and exclude any other unidentified latent entities. In this retrieval query, the placeholder for the target latent entity is replaced with the corresponding textual reference, as specified in the # Latent Entities section. All [SEP] tokens are removed.
The {infilling_query} is constructed by concatenating all relevant triplets from both the # Latent Entities and # Triples sections that include the target latent entity and exclude any other unidentified latent entities. In this case, the placeholder for the target latent entity is replaced with the special token <extra_id_0> to indicate the blank to be infilled. As above, all [SEP] tokens are removed.
See Table 7 for a detailed example of how the retrieval and infilling queries are constructed.
Consider the following graph: # Latent Entities:
(ENT1) [SEP] is [SEP] a musician
(ENT2) [SEP] is [SEP] a band
# Triples:
(ENT1) [SEP] is part of [SEP] Tall Birds
(ENT1) [SEP] is a percussionist for [SEP] (ENT2)
(ENT2) [SEP] formed in [SEP] Issaquah, Washington
In the case where (ENT1) is the first target latent entity to be infilled, the corresponding retrieval and infilling queries are: Retrieval Query: a musician is part of Tall Birds. Infilling Query: <extra_id_0> is part of Tall Birds. <extra_id_0> is a musician. After (ENT1) is identified as “Randall Nieman”, (ENT2) becomes the next target latent entity to be infilled. The corresponding retrieval and infilling queries are: Retrieval Query: Randall Nieman is a percussionist for a band. a band formed in Issaquah, Washington. Infilling Query: Randall Nieman is a percussionist for <extra_id_0>. <extra_id_0> formed in Issaquah, Washington. <extra_id_0> is a band.
Table 7: Example of retrieval and infilling queries for latent entity infilling.
B.3 Direct (Fact Triplet Verification)
The prompt in Table 8 is used in both Direct and the fact triplet verification step of GraphCheck to assess whether a claim is true or false, given the retrieved evidence.
Evidence: {evidence}
Claim: {claim}
Is the claim true or false?
Answer:
Table 8: Prompt used for Direct and fact triplet verification.
In the case of Direct, {claim} refers to the original input claim. In the fact triplet verification step of GraphCheck, it refers to a triplet converted into a natural language sentence (i.e., with [SEP] tokens removed). {evidence} denotes either (i) the concatenation of top-$k$ documents retrieved using the original claim or the converted triplet as the query, or (ii) each individual document within the top-$k$.
B.4 Strategy Selection
The prompt in Table 9 is used for strategy selection—deciding whether to use Direct or more systematic GraphCheck to verify a given claim. Specifically, it asks if the retrieved evidence documents contain sufficient information to support or refute the claim. If the answer is “yes” (sufficient), the claim is verified using Direct. If the answer is “no” (insufficient), the claim is verified using GraphCheck.
Evidence: {evidence}
Claim: {claim}
Does the evidence contain sufficient information to support or refute the claim? Yes or no?
Answer:
Table 9: Prompt used for strategy selection.
Here, {claim} refers to the original input claim. {evidence} is the concatenation of the top-$k$ documents retrieved using the original claim as the retrieval query.